October 9, 2019

From Ciswikidb
Jump to: navigation, search

CEBAF for 7 Months

Web gannt view (Internal): http://opsweb/gantt/2019_Fall/fall2019Run-chart.html
Task View (Internal): http://opsweb/gantt/2019_Fall/fall2019Run-tasks.html

CIS - Status and Strategy to Fall Operations

  • Oct 7 - 11 : Gun bake, leak check, lasers ready, laser retro, laser training, laser PQB start
  • Oct 14 - 18 : laser PQB continues, Re-assemble instrumentation, HV cables, coordinate with PSS/Ops for HV ON, conditioning?
  • Oct 21 - 25 : Make a photocathode, remove SF6 conditioning resistor, crane stuff out, clean up, start making beam w/ PQB checks
  • Oct 28 - Nov 1 : Injector setup and PQB checkout

What fits and what doesn't

  • Very little else fits
  • I don't think we have time for the Load chamber, little time
  • Daniel will coordinate removing the big HVPS oil tank

Small shed - lessons learned

The sure thing "tee" will be informative, help us find a path forward
We knew this was a test of Kr processing w/ a NEG coating => maybe we just learned the answer?
We worry about contamination ... during polishing, during venting, with NEGs, all worth considering, but really how different than usual?

CREX jobs

  • Laser Room
Camera measurements (post fiber change)
Spinning LP for A/B/C/D: http://opsweb.acc.jlab.org/CSUEApps/atlis/task/19669
Transition Time recheck
IA setting zero-finding
  • Beam to FC#1
Cathode Rotation: http://opsweb.acc.jlab.org/CSUEApps/atlis/task/18799
IA/RTP/RWHP settings for all 4 lasers: http://opsweb.acc.jlab.org/CSUEApps/atlis/task/19670
Pos Diff Feedback to FC1/FC2: http://opsweb.acc.jlab.org/CSUEApps/atlis/task/18793
  • Beam to either FC2/ILD (end of injector)
90Hz beam motion fix?
PQB Health check to FC2
PQB Health check to ILD
  • Beam to HallA
Adiabatic Damping check?
Dithering check?
  • During Experiments (affects other halls, but also necessary for HallC to ensure we are minimizing their Aq well)
PITA scans
RHWP sweeps
  • Post Experiment
Spot size asymmetry measurements: http://opsweb.acc.jlab.org/CSUEApps/atlis/task/19669

LERF

Nov 1-4 HCO for isotope run (laser, photocathode, high voltage)
Nov 5-6 : 2 shifts/day commissioning run, ~60 pC @ 4 MHz, < 1 mA
Dec TBD : another couple shifts, actual isotope run, ~60 pC @ 4 MHz < 1 mA

UITF

  • Electrode choice
  • Tentative dates for Booster commissioning
  • Wien filter testing
  • HD Ice run

GTS